OpenStudy (anonymous):

Evaluate the following expression using the values given. Find 4x − 2y − z, if x = 2, y = 3, and z = −8

OpenStudy (anonymous):

10

OpenStudy (anonymous):

4(2) - 2(3) - (-8) = 8 - 6 + 8 = 2 + 8 = 10

OpenStudy (anonymous):

0 8 −8 2

OpenStudy (anonymous):

10 is not a opition

OpenStudy (anonymous):

Then you may have copied the problem down wrong.

OpenStudy (anonymous):

its 0

OpenStudy (anonymous):

-4*-1-2*-2

OpenStudy (anonymous):

What??

OpenStudy (anonymous):

Look at the problem you wrote initially. It doesn't match up with what you just wrote.